# Nevanlinna Analytic Continuation for Migdal-Eliashberg Theory

### v1

@@ II.2 Using NAC with Migdal-Eliashberg Theory
verifying the correct mapping $\mathcal{N}:\mathbb{C}^{+}\rightarrow\overline{\mathbb{C}^{+}}$. The prove for $G^{22}_{n\mathbf{k}}(i\omega_{j})$ follows the same lines.
@@
$G^{\textrm{aux}}_{n\mathbf{k}}(i\omega_{j})=-\frac{2}{\Theta_{n\mathbf{k}}}\Big{[}\Big{.}$ $i\omega_{j}Z_{n\mathbf{k}}(i\omega_{j})+\Delta_{n\mathbf{k}}(i\omega_{j})Z_{n\mathbf{k}}(i\omega_{j})\Big{.}\Big{]}.$ (20)
@@
in agreement with Refs. [55, 53]. For the remaining complex functions in ME formalism, we have the following two equations
@@ III Results
We start by showing some results where NAC and the Padé approximation are in excellent agreement. First, for an isotropic full-bandwidth calculation of LaBeH8, with a fine $\mathbf{k}$-grid of $30\times 30\times 30$ at a temperature of $T=70$ K. We used default parameters for the Padé approximation as implemented in EPW. For NAC, we used the Pick criterion to determine the optimal number of Matsubara frequencies and a broadening parameter of $\eta=1$ meV. The same broadening was added to the spectral function from the Padé approximation in post-processing to obtain Fig. 3, where we show a comparison between the two methods for both the normal and anomalous spectral functions $A^{11}_{n\mathbf{k}}(\omega)$ and $A^{12}_{n\mathbf{k}}(\omega)$, respectively. The pole position, and therefore the quasiparticle energy renormalized by superconducting pairing, is located around $\omega=\pm 35$ meV. NAC is done without optimization, i.e. by using a constant free parameter of $\theta_{M}(z)=0$. We found that this works very well for the isotropic approximation to ME theory, although small unphysical satellite peaks with a maximum of $1$ eV-1 to $2$ eV-1 can appear at higher frequencies. The most common outcome is this kind of agreement between NAC and the Padé approximation, which we found to be the case in about 95% of our testing.
@@
Figure 5: Normal and anomalous spectral functions for LaBeH8 at a temperature of $T=130$ K within the isotropic full-bandwidth approximation obtained with EPW. Here NAC is able to avoid the numerical error of the Padé approximation around $\omega=\pm 10$ meV.
@@
Figure 6: Normal and anomalous spectral functions for MgB2 and LaBeH8 at a very low temperature of $T=0.05$ K within the isotropic Fermi-surface restricted approximation, obtained with EPW. The Padé approximation failed in these cases.
The numerical costs of NAC using quadruple precision are very similar to those of the Padé approximation using double precision. We reached this conclusion by comparing the runtimes of anisotropic calculations, as isotropic calculations are too brief to provide meaningful comparisons. However, it’s important to note that the runtimes of anisotropic calculations depend highly on the chosen $\mathbf{k}$-grid, so the following estimates are intended to give a general sense of the performance. In our calculations for MgB2, we observed typical CPU runtimes of approximately 50 seconds per temperature for NAC with quadruple precision, comparable to the runtimes for Padé approximation calculations using double precision. In contrast, using quadruple precision for the Padé approximation significantly increases the computation time to around 500 seconds.
@@ IV Conclusion
In this work, we demonstrate a method to obtain real-frequency properties from the analytically continued Nambu-Gor’kov matrix within the framework of Migdal-Eliashberg (ME) theory for superconductivity. By leveraging the symmetry properties of the Green’s function, our method requires only two analytic continuations to obtain all components in real frequency space: one for the normal and one for the auxiliary Green’s function. This significantly broadens the scope for applying a variety of analytic continuation techniques that were previously inaccessible in the Migdal-Eliashberg framework, retaining full information of the real frequency complex functions explicitely, such as $\Delta_{n\mathbf{k}}(\omega)$, $Z_{n\mathbf{k}}(\omega)$, and $\chi_{n\mathbf{k}}(\omega)$. Our generalized workflow opens up the possibility for employing advanced methods like MaxEnt, stochastic analytic continuation, genetic algorithms, and causal projection in the modeling of superconducting materials, thereby enhancing the applicability and accuracy of the Migdal-Eliashberg formalism.
@@ Appendix A Computational Details
After that, we used the Wannier implementation of EPW to interpolate electron-phonon matrix elements onto dens $\mathbf{k}$- and $\mathbf{q}$-grids. We used the ”selected columns of the density matrix” (SCDM) method [60, 61] to automatically wannierize our electronic band structure, using a frozen window of about [14[-14[ - 14 eV, $+5$ eV] around the Fermi energy $E_{F}$. For the SCDM entanglement, we use an error function with the two parameters $\sigma=5$ and $\mu_{c}$ at the Fermi energy $E_{F}$. Then for the solution of the isotropic full-bandwidth ME equations, we used a coarse $\mathbf{k}$- and $\mathbf{q}$-grids of $6\times 6\times 6$ interpolated onto $30\times 30\times 30$ fine grids. We set the lower boundary of phonon frequency to $10$ cm-1, the width of the Fermi surface window to 1 eV, and the imaginary convergence threshold to $10^{-5}$ Ry, all for a Morel-Anderson pseudopotential of $\mu^{*}=0.16$. The ME equations were solved while updating the chemical potential to keep the particle number constant, which requires a comparably high-frequency cut-off of $10$ eV for convergence. We used sparse sampling and analytically continued with the Padé approximation and NAC simultaneously, starting at a temperature of $60$ K increased by steps of $10$ K.

### v2

@@ II.2 Using NAC with Migdal-Eliashberg Theory
verifying the correct mapping $\mathcal{N}:\mathbb{C}^{+}\rightarrow\overline{\mathbb{C}^{+}}$ [6, 7]. The prove for $G^{22}_{n\mathbf{k}}(i\omega_{j})$ follows the same lines.
@@
$G^{\textrm{aux}}_{n\mathbf{k}}(i\omega_{j})=-\frac{2}{\Theta_{n\mathbf{k}}(i\omega_{j})}\Big{[}\Big{.}$ $i\omega_{j}Z_{n\mathbf{k}}(i\omega_{j})+\Delta_{n\mathbf{k}}(i\omega_{j})Z_{n\mathbf{k}}(i\omega_{j})\Big{.}\Big{]}.$ (20)
@@
in agreement with Refs. [55, 54, 53]. For the remaining complex functions in ME formalism, we have the following two equations
@@ III Results
We start by showing some results where NAC and the Padé approximation are in excellent agreement. First, for an isotropic full-bandwidth calculation of LaBeH8, with a fine $\mathbf{k}$-grid of $30\times 30\times 30$ at a temperature of $T=70$ K. We used default parameters for the Padé approximation as implemented in EPW. For NAC, we used the Pick criterion to determine the optimal number of Matsubara frequencies and a broadening parameter of $\eta=1$ meV. The same broadening was added to the spectral function from the Padé approximation in post-processing to obtain Fig. 3, where we show a comparison between the two methods for both the normal and anomalous spectral functions $A^{11}_{n\mathbf{k}}(\omega)$ and $A^{\textrm{an}}_{n\mathbf{k}}(\omega)$, respectively. The pole position, and therefore the quasiparticle energy renormalized by superconducting pairing, is located around $\omega=\pm 35$ meV. NAC is done without optimization, i.e. by using a constant free parameter of $\theta_{M}(z)=0$. We found that this works very well for the isotropic approximation to ME theory, although small unphysical satellite peaks with a maximum of $1$ eV-1 to $2$ eV-1 can appear at higher frequencies. The most common outcome is this kind of agreement between NAC and the Padé approximation, which we found to be the case in about 95% of our testing.
@@
Figure 5: Normal and anomalous spectral functions for LaBeH8 at a temperature of $T=130$ K within the isotropic full-bandwidth approximation obtained with EPW. Here, NAC is able to avoid the numerical error of the Padé approximation at around $\omega=\pm 110$ meV.
@@
Figure 6: Normal spectral functions for MgB2 and LaBeH8 at a very low temperature of $T=0.05$ K within the isotropic Fermi-surface restricted approximation, obtained with EPW. The Padé approximation failed in these cases.
The numerical costs of NAC using quadruple precision are very similar to those of the Padé approximation using double precision. We reached this conclusion by comparing the runtimes of anisotropic calculations, as isotropic calculations are too brief to provide meaningful comparisons. However, it’s important to note that the runtimes of anisotropic calculations depend highly on the chosen $\mathbf{k}$-grid, so the following estimates are intended to give a general sense of the performance. In our calculations for MgB2, we observed typical CPU runtimes of approximately $50$s per temperature for NAC with quadruple precision, comparable to the runtimes for Padé approximation calculations using double precision. In contrast, using quadruple precision for the Padé approximation significantly increases the computation time to around $500$s.
@@ IV Conclusion
In this work, we demonstrate a method to obtain real-frequency properties from the analytically continued Nambu-Gor’kov matrix within the framework of Migdal-Eliashberg (ME) theory for superconductivity. By leveraging the symmetry properties of the Green’s function, our method requires only two analytic continuations to obtain all components in real frequency space: one for the normal and one for the auxiliary Green’s function. This significantly broadens the scope for applying a variety of analytic continuation techniques that were previously inaccessible in the Migdal-Eliashberg framework, retaining full information of the real frequency complex functions explicitly, such as $\Delta_{n\mathbf{k}}(\omega)$, $Z_{n\mathbf{k}}(\omega)$, and $\chi_{n\mathbf{k}}(\omega)$. Our generalized workflow opens up the possibility for employing advanced methods like MaxEnt, stochastic analytic continuation, genetic algorithms, and causal projection in the modeling of superconducting materials, thereby enhancing the applicability and accuracy of the Migdal-Eliashberg formalism.
@@ Appendix A Computational Details
After that, we used the Wannier implementation of EPW to interpolate electron-phonon matrix elements onto dense $\mathbf{k}$- and $\mathbf{q}$-grids. We used the ”selected columns of the density matrix” (SCDM) method [60, 61] to automatically wannierize our electronic band structure, using a frozen window of about [14[-14[ - 14 eV, $+5$ eV] around the Fermi energy $E_{F}$. For the SCDM entanglement, we use an error function with the two parameters $\sigma=5$ and $\mu_{c}$ at the Fermi energy $E_{F}$. Then for the solution of the isotropic full-bandwidth ME equations, we used a coarse $\mathbf{k}$- and $\mathbf{q}$-grids of $6\times 6\times 6$ interpolated onto $30\times 30\times 30$ fine grids. We set the lower boundary of phonon frequency to $10$ cm-1, the width of the Fermi-surface window to 1 eV, and the imaginary convergence threshold to $10^{-5}$ Ry, all for a Morel-Anderson pseudopotential of $\mu^{*}=0.16$. The ME equations were solved while updating the chemical potential to keep the particle number constant, which requires a comparably high frequency cut-off of $10$ eV for convergence. We used sparse sampling and analytically continued with the Padé approximation and NAC simultaneously, starting at a temperature of $60$ K increased by steps of $10$ K.